# Enhancing NMR Shielding Predictions of Atoms-in-Molecules Machine Learning Models with Neighborhood-Informed Representations

### v1

@@ IV.2 Transferability to external datasets of larger molecules
To evaluate transferability, we first test the descriptors on larger drug-like molecules from the Drug12 and Drug40 datasets. As summarized in Table 3, the predicted MAEs with aBoB-RBF(4) are lower than those obtained previously with FCHL Gupta, Chakraborty, and Ramakrishnan (2021) and MACE-OFF23-small Shiota, Ishihara, and Mizukami (2024), showing that this descriptor can capture chemical environments not only for small QM9 molecules but also for larger systems. For the QM9NMR dataset, MAEs are significantly larger than the nearly vanishing mean signed errors (MSEs), indicating that the deviations from the DFT reference values are largely non-systematic and cancel out between positive and negative errors. By contrast, for Drug40 molecules (7–17 CONF atoms), the aBoB-RBF(4) descriptor exhibits slightly negative MSEs, suggesting a mild underestimation of chemical shifts as system size increases. This bias becomes more pronounced in Drug12 molecules (17–23 CONF atoms), where the MSE reaches about 1-1 ppm. Correcting for this systematic bias with a few reference values would lower the MAE to approximately 2.4 ppm. For comparison, the aSLATM(2) descriptor yields consistently small MSEs, reflecting a lower degree of systematic bias; however, its residual prediction errors remain non-systematic and thus less amenable to post-prediction calibration.
@@
Finally, we test 200 molecules from the GDBm dataset, a subset of the GDB17 database Ruddigkeit et al. (2012). As shown in Figure 5b, predictions of 13C $\sigma_{\rm iso}$ values yield $R^{2}=0.995$ and an MAE of 2.66 ppm. As the number of heavy atoms $m$ increases, MAEs gradually rise (up to 3.06 ppm for $m=16$), and while $R^{2}$ remains at 0.99, consistent with the fact that the models were trained on molecules containing fewer CONF atoms.
@@ IV.3 Example NMR prediction using the mlqm9nmr module
The 100k models developed in this study are distributed through an open-source Python package, mlqm9nmr mlq (2025), which enables prediction of 13C chemical shifts for out-of-sample molecules. A usage example with the aBoB-RBF($4$) descriptor is shown in Figure 6. Installation and usage instructions are provided in the SI. The package includes trained models for aCM(4), aCM-RBF(4), aBoB(4), and aBoB-RBF(4).
@@ V Conclusions
In this study, we have demonstrated that incorporating information from the nearest neighbors enhances the accuracy of the descriptors. With the augmentation of the descriptors of $n$ neighbors to the parent atom’s descriptor, the improvement in the accuracy of the $^{13}C$ chemical shift is gradual, but it saturates for $n=5$. As the maximum number of bonded neighbors for a C-atom is 4 ($sp^{3}$ hybridized), the choice of $n=4$ offers the least prediction errors.
@@
To demonstrate transferability, the models were further validated on out-of-sample datasets, including Drug12, Drug40, pyrimidinone, and GDBm, achieving MAEs of 3.43, 2.80, 2.19, and 2.66 ppm, respectively. These results show that the ML models developed in this study are not only accurate for small QM9 molecules but also transferable to larger and more diverse chemical systems. Additionally, the $\Delta$-ML approach Ramakrishnan et al. (2015), not explored here, offers a promising route to enhance accuracy further and compensate for differences between semi-empirical and DFT-level geometries, thereby pushing performance closer to high-level theoretical benchmarks. It has not escaped our attention that the aBoB-RBF(4) descriptor, while demonstrated here for NMR shieldings, may find broader utility in modeling other atom-specific properties such as core-electron binding energies and atomic partial charges.

### v3

@@ IV.2 Transferability to external datasets of larger molecules
To evaluate transferability, we first test the descriptors on larger drug-like molecules from the Drug12 and Drug40 datasets. As summarized in Table 3, the predicted MAEs with aBoB-RBF(4) are lower than those obtained previously with FCHL gupta2021revving and MACE-OFF23-small shiota2024universal, showing that this descriptor can capture chemical environments not only for small QM9 molecules but also for larger systems. For the QM9NMR dataset, MAEs are significantly larger than the nearly vanishing mean signed errors (MSEs), indicating that the deviations from the DFT reference values are largely non-systematic and cancel out between positive and negative errors. By contrast, for Drug40 molecules (7–17 CONF atoms), the aBoB-RBF(4) descriptor exhibits slightly negative MSEs, suggesting a mild underestimation of chemical shifts as system size increases. This bias becomes more pronounced in Drug12 molecules (17–23 CONF atoms), where the MSE reaches about 1-1 ppm. Correcting for this systematic bias with a few reference values would lower the MAE to approximately 2.4 ppm. For comparison, the aSLATM(2) descriptor yields consistently small MSEs, reflecting a lower degree of systematic bias; however, its residual prediction errors remain non-systematic and thus less amenable to post-prediction calibration.
@@
Finally, we test 200 molecules from the GDBm dataset, a subset of the GDB17 database ruddigkeit2012enumeration. As shown in Figure 5b, predictions of 13C $\sigma_{\rm iso}$ values yield $R^{2}=0.995$ and an MAE of 2.66 ppm. As the number of heavy atoms $m$ increases, MAEs gradually rise (up to 3.06 ppm for $m=16$), and while $R^{2}$ remains at 0.99, consistent with the fact that the models were trained on molecules containing fewer CONF atoms.
@@ IV.3 Example NMR prediction using the mlqm9nmr module
The 100k models developed in this study are distributed through an open-source Python package, mlqm9nmr mlqm9nmr, which enables prediction of 13C chemical shifts for out-of-sample molecules. A usage example with the aBoB-RBF($4$) descriptor is shown in Figure 6. Installation and usage instructions are provided in the SI. The package includes trained models for aCM(4), aCM-RBF(4), aBoB(4), and aBoB-RBF(4).
@@ V Conclusions
In this study, we have demonstrated that incorporating information from the nearest neighbors enhances the accuracy of the descriptors. With the augmentation of the descriptors of $n$ neighbors to the parent atom’s descriptor, the improvement in the accuracy of the 13C chemical shift is gradual, but it saturates for $n=5$. As the maximum number of bonded neighbors for a C-atom is 4 ($sp^{3}$ hybridized), the choice of $n=4$ offers the least prediction errors.
@@
To demonstrate transferability, the models were further validated on out-of-sample datasets, including Drug12, Drug40, pyrimidinone, and GDBm, achieving MAEs of 3.43, 2.80, 2.19, and 2.66 ppm, respectively. These results show that the ML models developed in this study are not only accurate for small QM9 molecules but also transferable to larger and more diverse chemical systems. Using the $\Delta$-ML approach ramakrishnan2015big with the aBoB-RBF(4) descriptor, the MAE for predicting 50k hold-out 13C values from the QM9NMR dataset at the target level mPW1PW91/6-311+G(2d,p) drops to 1.16 ppm when trained on 100k values and using mPW1PW91/STO-3G as the baseline. Comparable errors were also observed when HF/STO-3G was used as the baseline. This represents a substantial improvement over direct prediction of the target level (MAE, 1.69 ppm), bringing the error closer to that required for structure elucidation ($\approx 1$ ppm). Hence the use of $\Delta$-ML offers a promising route to enhance accuracy further and compensate for differences between semi-empirical and DFT-level geometries, thereby pushing performance closer to high-level theoretical benchmarks. It has not escaped our attention that the aBoB-RBF(4) descriptor, while demonstrated here for NMR shieldings, may find broader utility in modeling other atom-specific properties such as core-electron binding energies and atomic partial charges.